OpenStudy (anonymous):

The height of an equilateral triangular prism increases by three units. The new volume is more than the original by how much?

OpenStudy (anonymous):

example for A = 1/2 *3*4*3 (this is height) = 18 for B = 1/2*3*4*6(increase 3 units)= 36 the volume will increase two times from the original